# Harnessing The Collective Wisdom: Fusion Learning Using Decision Sequences From Diverse Sources

### v2

@@ Abstract
Learning from the collective wisdom of crowds is related to the statistical notion of fusion learning from multiple data sources or studies. However, fusing inferences from diverse sources is challenging since cross-source heterogeneity and potential data-sharing complicate statistical inference. Moreover, studies may rely on disparate designs, employ myriad modeling techniques, and prevailing data privacy norms may forbid sharing even summary statistics across the studies for an overall analysis. We propose an Integrative Ranking and Thresholding (IRT) framework for fusion learning in multiple testing. IRT operates under the setting where from each study a triplet is available: the vector of binary accept-reject decisions on the tested hypotheses, its False Discovery Rate (FDR) level and the hypotheses tested by it. Under this setting, IRT constructs an aggregated and nonparametric measure of evidence against each null hypotheses, which facilitates ranking the hypotheses in the order of their likelihood of being rejected. We show that IRT guarantees an overall FDR control if the studies control their respective FDR at the desired levels. IRT is extremely flexible, and a comprehensive numerical study demonstrates its practical relevance for pooling inferences. A real data illustration and extensions to alternative forms of Type I error control are discussed.
Keywords: Crowdsourcing, E-values, False Discovery Rate, Integrative inference, Meta-analysis
@@ 1 Introduction
Learning from the wisdom of crowds is the process of synthesizing the collective wisdom of disparate participants for performing related tasks, and has evolved into an indispensable tool for modern scientific inquiry. This process is also known as ‘Crowdsourcing’, which has generated tremendous societal benefits through novel drug discovery (Chodera et al., 2020), new machine learning algorithms (Sun et al., 2022), product design (Jiao et al., 2021) and crime control (Logan, 2020), to name a few. Harnessing such collective wisdom allows incorporating diverse opinions as well as expertise into the decision-making process and ultimately broadens the transparency of scientific findings (Surowiecki, 2005).
Synthesizing the collective wisdom of crowds is related to the statistical notion of fusion learning from multiple data sources or studies (Liu et al., 2022; Guo et al., 2023). Over the past decade, techniques for such fusion learning have found widespread use for incorporating heterogeneity into the underlying analysis and increasing the power of statistical inference. For instance, in recent years the accrescent volume of gene expression data available in public data repositories, such as the Gene Expression Omnibus (GEO) (Barrett et al., 2012) and Array Express (Rustici et al., 2012), have facilitated the synthesis and reuse of such data for meta-analysis across multiple studies. Biologists can use OMiCC (Shah et al., 2016), a crowdsourcing web platform, for generating and testing hypotheses by integrating data from diverse studies. Relatedly, sPLINK (Nasirigerdeh et al., 2022), a hybrid federated tool, allows conducting privacy-aware genome-wide association studies (GWAS) on distributed datasets. In distributed multisensor systems, such as wireless sensor networks, each node in the network evaluates its designated region in space by conducting simultaneous test of several hypotheses. The decision sequence is then transmitted to a fusion center for processing. Sensor fusion techniques are used to synthesize the data from multiple sensors, which provides improved accuracy and statistically powerful inference than that achieved by the use of a single sensor alone (Hall and Llinas, 1997; Shen and Wang, 2001; Jamoos and Abuawwad, 2020).
However, fusing inferences from diverse sources [^1] We will use the terms ‘data-source’, ‘study’ and ‘nodes’ interchangeably throughout this article. is challenging for several reasons. First, cross-source heterogeneity and potential data-sharing complicate statistical inference. For instance, in microarray meta-analysis, an integrative false discovery rate (FDR) analysis of the multiple hypotheses often requires making additional assumptions such as study independence and strong modeling assumptions to capture between-study heterogeneity. Second, prevailing data privacy norms may forbid sharing even summary statistics across the studies for an overall analysis. This is particularly relevant in the context of genomic data where NIH (National Institutes of Health) restricts the availability of dbGaP (database of Genotypes and Phenotypes) data to approved users (Couzin, 2008; Zerhouni and Nabel, 2008). Also, in wireless sensor networks, communication limits may restrict sharing all information from the nodes to the fusion center. Third, in the case of meta-analysis where fusion learning is widespread, different studies may rely on disparate designs and widely different modeling techniques for individual inferences. Besides introducing algorithmic randomness into the individual inferences, it is also unclear how such inferences can be integrated and subsequently interpreted. Fourth, often significant statistical and computational experience is required for integrating the individual inferences. This may preclude investigators without substantial statistical training to perform such analyses.
@@
In this paper, we develop a framework for fusion learning in multiple testing that seeks to overcome the aforementioned challenges of integrative inference across multiple data sources. Our framework, which we call IRT for Integrative Ranking and Thresholding, operates under the setting where from each study a triplet is available: the study-specific vector of binary accept / reject decisions on the tested hypotheses, the FDR level of the study and the hypotheses tested by the study. Under this setting, the IRT framework consists of two key steps: in step (1) IRT utilizes the binary decisions from each study to construct nonparametric evidence indices which serve as measures of evidence against the corresponding null hypotheses, and in step (2) the evidence indices across the studies are fused into a discriminatory measure that ranks the hypotheses in the order of their likelihood of being rejected. The proposed fusion learning framework has several distinct advantages. First, the IRT framework guarantees an overall FDR control as long as the individual studies control the FDR at their desired levels. This FDR control holds under arbitrary dependence between the fused evidence indices from step (2). Second, IRT is extremely simple to implement and is broadly applicable without any model assumptions. This particular aspect is especially appealing because IRT synthesizes inferences from diverse studies irrespective of the underlying multiple testing algorithms employed by the studies. Third, the evidence indices in our framework are closely related to “$e-$values” (see Shafer (2021); Vovk and Wang (2021); Grünwald et al. (2023); Ramdas et al. (2020); Wasserman et al. (2020) for an incomplete list) for hypothesis testing. Besides being a natural counterpart to the popular $p-$values in statistical inference, $e-$values are relatively more robust to model misspecification and particularly to dependence between the $p-$values. In our numerical experiments we find that when the $p-$values are dependent, IRT is substantially more powerful than methods that rely on a conversion from $p-$values to $e-$values for pooling inferences. Additionally, when the distribution of the test statistic under the null hypothesis is misspecified and the corresponding $p-$values are sub-optimal, procedures that rely directly on the magnitude of such $p-$values for pooling inferences may fail to control the FDR. In contrast, IRT is relatively robust to such misspecification in our simulation studies. For almost a century, $e-$values have been used in Statistics, often disguised as likelihood ratios, Bayes factors, stopped nonnegative supermartingales. See Ramdas et al. (2022) for a survey on additional settings, such as betting scores, game-theoretic statistics and safe anytime-valid inference, where $e-$values arise naturally. Finally, to the best of our knowledge, IRT is the first fusion learning framework for multiple testing that relies on the binary decision sequences which are relatively more private than summary statistics. We rely on a novel construction of $e-$values from the accept-reject decisions which are also related to the all-or-nothing bet of Shafer (2021) for testing a single null hypothesis.
Figure 1: A pictorial representation of the IRT framework. In panel (1) $d$ studies are testing a study-specific subset of $m$ null hypotheses. Here the black solid squares represent hypotheses that are not tested by the individual studies. In panel (2), these studies use different multiple testing algorithms that are designed to control their respective FDR at level $\alpha_{j},~{}j=1,\ldots,d$. The ‘X’-marked squares in panel (3) depict the rejected hypotheses and represent the data that is available to the IRT framework. It includes the binary decision vectors $\delta_{ij}$ from each study, the study-sepcific FDR level $\alpha_{j}$ and the subset of $m$ hypotheses evaluated by each study. In panel (4), the IRT framework converts each binary decision into a nonparametric evidence index such that a higher index, displayed in a darker shade of red, represents a larger likelihood for rejecting that study-specific null hypothesis. In panel (5), the study-specific evidence indices are fused in a manner such that the aggregated indices rank the $m$ hypotheses and, as demonstrated in panel (6), the rankings can be used to determine a cutoff for valid FDR control at a pre-determined level $\alpha$.
1.1 Outline of the proposed approach
Figure 1 provides a pictorial representation of the IRT framework under a simplified setting. A precise formulation and formal mathematical statements are deferred until Section 2. Panel (1) of Figure 1 represents a scenario where $d$ studies are testing a study-specific subset of $m$ null hypotheses. Here the black solid squares represent hypotheses that are not tested by the individual studies. In Panel (2), these individual studies use different multiple testing algorithms that are designed to control their respective FDR at level $\alpha_{j},~{}j=1,\ldots,d$. The ‘X’-marked squares in Panel (3) depict the rejected hypotheses and represent the data that is available to the IRT framework for synthesis. It includes the binary decisions $\delta_{ij}$ from each study, the study-specific FDR level $\alpha_{j}$ and the subset of $m$ hypotheses evaluated by each study. Panels (4)-(6) illustrate the IRT framework. Particularly, in Panel (4), the IRT framework converts each binary decision $\delta_{ij}$ into a nonparametric evidence index such that a higher index, displayed in a darker shade of red, represents a stronger conviction against that study-specific null hypothesis. In Panel (5), the study-specific evidence indices are fused in a manner such that the aggregated indices rank the $m$ hypotheses and, as demonstrated in Panel (6), the rankings can be used to determine a cutoff for valid FDR control at a pre-determined level $\alpha$. While Figure 1 introduces IRT under a setting where $d$ studies are testing a study-specific subset of $m$ null hypotheses, it is much more broadly applicable, including the setting where pooling inferences for a common set of $m$ hypotheses across the $d$ studies is of interest. Section 2 presents those details.
1.2 Related literature and main methodological contributions
The IRT framework is related to multiple strands of literature, which includes FDR control using $e-$values, integrative multiple testing under data-privacy and algorithmic derandomization. We discuss how IRT differs from these existing body of works.
Multiple testing with $e-$values - Recently, there has been a proliferation of interest in developing methods for simultaneous hypotheses testing using $e$-values. See, for instance, Wang and Ramdas (2022); Ignatiadis et al. (2024); Chi et al. (2022); Vovk and Wang (2023); Xu and Ramdas (2023) and the references therein. In these works, the $e-$values are typically constructed from either $p-$values or likelihood ratios. In contrast, the generalized $e-$values (Wang and Ramdas, 2022; Ren and Barber, 2024) in our setting arise from the accept / reject decisions of the corresponding null hypotheses and are constructed in a nonparametric fashion.
Integrative multiple testing under data-privacy - The literature on the evolving field of integrative multiple testing under data-privacy is also related to our work. For instance, under the restricted data-sharing mechanism of Wolfson et al. (2010), Liu et al. (2021) propose an integrative high-dimensional multiple testing framework with asymptotic FDR control. Our work differs from Liu et al. (2021) in two main aspects. First, we consider a relatively more stringent data-privacy regime where the studies are only allowed to share their decision sequences, the FDR level and the hypotheses tested by them. Second, the integrative FDR control offered by IRT is non-asymptotic if the individual studies guarantee a non-asymptotic control of their respective FDR levels (see Section 4.2). This aspect is particularly appealing when the underlying multiple testing problem is small-scale.
Algorithmic stability - Numerous methods have been proposed for mitigating algorithmic randomness in recent years. For instance, Bashari et al. (2023) propose a method to aggregate conformal tests for outliers obtained with repeated splits of the same data set, while controlling the FDR. In the context of high dimensional variable selection, Ren and Barber (2024) develop a methodology for derandomizing model-X knockoffs (Barber and Candès, 2015; Candes et al., 2018) with provable FDR control while Dai et al. (2023a, b) develop a multiple data-splitting method with asymptotic FDR control to stabilize variable selection results. The IRT framework can be similarly viewed as a method for mitigating algorithmic randomness that may potentially arise from the use of myriad multiples testing procedures for FDR control by the $d$ studies. Crucially, however, our framework does not rely on hard-to-verify assumptions for FDR control, such as Dai et al. (2023a, b), and in contrast to Ren and Barber (2024); Bashari et al. (2023); Li and Zhang (2023), IRT only requires access to the binary decision vector of each study for integrative FDR control. In fact, in Section 4.1 we show that several recent works that leverage $e-$values for aggregating and derandomizing specific FDR methods (see Li and Zhang, 2023; Ren and Barber, 2024; Zhao and Sun, 2024) are special cases of our evidence index (Equation (1)) from an asymptotic perspective.
The ensuing discussion summarizes our main methodological contributions.
  1. 1.
    Construction of evidence indices from binary decisions - at an intuitive level, a multiple testing rule essentially translates evidence indices, such as $p-$values, $e-$values, $z-$scores, into binary decisions. This paper presents a comprehensive solution to the “reverse testing problem”, which involves converting binary decisions into evidence indices. We choose $e-$value as an evidence index because of its many desirable theoretical properties as pointed out in Wang and Ramdas (2022); Vovk and Wang (2021). While the conversion of binary decisions to $e-$values has been explored in Ren and Barber (2024); Li and Zhang (2023), these studies are limited to specific testing processes, such as knockoffs and BH, and focus solely on testing rules that guarantee FDR control. In contrast, we show how $e-$values can be constructed from any testing procedure that guarantees some kind of type I error control, such as FDR, $k-$FWER, and PFER (see Section 7).
  2. 2.
    Evidence aggregation via multiplication - for each hypothesis, IRT fuses the evidence indices via their arithmetic mean. However, when inferences are pooled with the goal of higher reliability then such an aggregation scheme often renders IRT extremely conservative. In Section 3.2, we provide precise conditions under which the evidence indices can be fused using a product rule. These conditions are related to the notion of partial exchangeablilty on the set of nulls, commonly used in the conformal inference literature (Bates et al., 2023; Liang et al., 2024), and symmetric decision rules, which arise naturally in conventional settings where all hypotheses undergo simultaneous testing without the inclusion of auxiliary side information. The underlying inference pooling framework that relies on this product aggregation scheme is called IRT and is, typically, more powerful than IRT.
  3. 3.
    A flexible framework for pooling inferences - while the proposed framework is extremely simple, it enjoys substantial flexibility with which it can be used in myriad applications. For instance, IRT(IRT) is applicable when only some of the studies share $p-$values, while for the others only the binary decisions are available. In such a setting, IRT (IRT) can be used in conjunction with a p2e calibrator, such as those proposed in Vovk and Wang (2021), for pooling inferences (see Remark 3). Furthermore, IRT (IRT) can pool inferences across studies for an overall FDR control even when studies employ different type I error control metrics, such as family-wise error rate or per-family error rate control (see Remark 4). Finally, in many practical settings prior knowledge may dictate that some studies share data amongst themselves or rely on auxiliary side information for multiple testing. In this scenario, a hybrid scheme IRT H, that combines IRT and IRT, can synthesize inferences across all the studies and ensure valid FDR control (see Remark 1).
  4. 4.
    Robustness to misspecified null distribution - when the summary statistics, such as $p-$values, are available from all studies then neither IRT nor IRT are natural candidates for pooling inferences as they may lead to power loss. This is indeed true as long as the null distribution of the test statistic is correctly specified. The numerical experiments in Section 6 reveal that when the null distribution is misspecified, methods that rely directly on the magnitude of the $p-$values, may fail to control FDR at the desired level. In contrast IRT, IRT and IRT H are relatively robust to such misspecification. This is an important practical setting since typically the null distribution specification relies on large sample approximations, and such large samples may not be available for some studies.
1.3 Organization
The article is laid out as follows: Section 2 presents the IRT framework for integrative FDR control. In Section 3 we present IRT that relies on an alternative scheme for evidence aggregation. Additional insights about the IRT framework are discussed in Section 4. A real data illustration is presented in Section 5 while Section 6 reports the empirical performance of IRT on synthetic data. Extensions to alternative forms of Type I error control are discussed in Section 7. The article concludes with a summary in Section 8. The data and R-code used in this article are available at https://github.com/trambakbanerjee/IRT.
Notations
Let =$\mathcal{M}=\{1,\dots,m\}$ and consider testing $m$ null hypotheses $H_{01},\dots,H_{0m}$. Denote $\mathcal{H}_{0}=\{i:H_{0i}\ \text{is true}\}$ and $\mathcal{H}_{1}=\mathcal{M}\setminus\mathcal{H}_{0}$, respectively, as the set of true null and non-null hypotheses. Here $\setminus$” denotes the usual set difference operator. Let $\mathbb{I}(\cdot)$ be the indicator function that returns 1 if the condition is met and 0 otherwise, and denote $\|\mathbf{w}\|_{p}$ as the $\ell_{p}$-norm of vector $\mathbf{w}$. In the sequel, $a\vee b$ will denote $\max\{a,b\}$ for two real numbers $a$ and $b$, $[d]=\{1,\ldots,d\}$ and the cardinality of a set of positive integers $\mathcal{I}$ will be denoted by $|\mathcal{I}|$.
@@ 2 IRT: integrative FDR control using binary decision sequences
We consider a setting where the $\mathcal{M}$ hypotheses are tested by $d$ studies with study $j$ testing $\mathcal{M}_{j}$ hypotheses. Here $\mathcal{M}_{j}\cap\mathcal{M}\neq\emptyset$ and $|\mathcal{M}_{j}|=m_{j}$ for all $j\in[d]$. Let $\theta_{i}=\mathbb{I}(H_{0i}\text{~{}is false})$ be an indicator function that gives the true state of the $i$-th testing problem and denote $\delta_{ij}\in\{0,1\}$ as the decision that study $j$ makes about hypothesis test $i\in\mathcal{M}_{j}$, with $\delta_{ij}=1$ being a decision to reject $H_{0i}$. Denote the vector of decisions as $\bm{\delta}_{j}=(\delta_{1j},\cdots,\delta_{m_{j}j})\in\{0,1\}^{m_{j}}$. A selection error, or false positive, occurs if study $j$ asserts that $H_{0i}$ is false when it actually is not. In multiple testing problems, such false positive decisions are inevitable if we wish to discover interesting effects with a reasonable power. Instead of aiming to avoid any false positives, a practical goal is to keep the false discovery rate (FDR, Benjamini and Hochberg, 1995) small, which is the expected proportion of false positives among all selections,
$\text{FDR}(\bm{\delta}_{j})=\mathbb{E}\left[\text{FDP}(\bm{\delta}_{j})\right]~{}\text{where FDP}(\bm{\delta}_{j})=\dfrac{\sum_{i\in\mathcal{M}_{j}}(1-\theta_{i})\delta_{ij}}{\max\Bigr{\{}\sum_{i\in\mathcal{M}_{j}}\delta_{ij},1\Bigr{\}}}.$
The power of a testing procedure is measured by the expected number of true positives (ETP) where,
$\mbox{ETP}(\bm{\delta}_{j})=\mathbb{E}\left[\sum_{i\in\mathcal{M}_{j}}\theta_{i}\delta_{ij}\right]$
Hence, the multiple testing problem for study $j$ can be formulated as
$\text{maximize}_{\bm{\delta}_{j}}~{}\text{ETP}(\bm{\delta}_{j})~{}\text{subject to }\mbox{FDR}(\bm{\delta}_{j})\leq\alpha_{j},$
where $\alpha_{j}\in(0,1)$ is a pre-specified cap on the maximum acceptable FDR for the $j$-th study. Our goal is to conduct inference for $\bm{\theta}=(\theta_{1},\ldots,\theta_{m})$ by fusing the evidence available in the binary decision sequences $\bm{\delta}_{j}$ provided by the $d$ studies. We develop an integrative ranking and thresholding procedure that operates on the triplet $\mathcal{D}_{j}=\{\bm{\delta}_{j},\alpha_{j},\mathcal{M}_{j}\}$ available from each study and provably controls the FDR within level $\alpha\in(0,1)$.

### v3

@@ Abstract
We introduce an Integrative Ranking and Thresholding (IRT) framework for fusing evidence from multiple testing procedures. The key innovation is a method that transforms binary testing decisions into compound $e$-values, enabling the combination of findings across diverse data sources or studies. We demonstrate that IRT ensures overall false discovery rate (FDR) control, provided the individual studies maintain their respective FDR levels. This approach is highly flexible and is a powerful alternative for fusing inferences in meta-analysis where some studies report summary statistics while the rest reveal only the rejections under a pre-specified FDR level. Extensions to alternative Type I error control measures are explored.
Keywords: E-values; False Discovery Rate; Integrative inference; Meta-analysis.
@@ 1 Introduction
Synthesizing the collective wisdom of crowds is related to the statistical notion of fusion learning. However, fusing inferences from diverse sources [^1] We will use the terms ‘data-source’ and ‘study’ interchangeably throughout this article. is challenging for several reasons. First, cross-source heterogeneity and potential data-sharing complicate statistical inference, often requiring strong assumptions like study independence. Second, many existing meta-analytic tools require continuous summary statistics, such as $p$-values. However, it is common for some studies to only report a binary list of discoveries from an FDR-controlled procedure (Tang et al., 2014). Under study dependence, contemporary methods are unable to coherently integrate these mixed-evidence formats. Third, disparate experimental designs and modeling techniques yield outputs that are not directly comparable, posing a significant hurdle towards their integration. Fourth, performing such integrative analyses often requires specialized statistical expertise, limiting their broader application.
@@
In this work, we propose a general and flexible framework for fusing multiple statistical testing decisions, which we call IRT for Integrative Ranking and Thresholding. IRT operates under the setting where from each study a triplet is available: the study-specific vector of binary accept / reject decisions on the tested hypotheses, the FDR level of the study and the hypotheses tested by the study. Under this setting, the IRT framework consists of two key steps: in step (1) IRT utilizes the binary decisions from each study to construct nonparametric evidence indices which serve as measures of evidence against the corresponding null hypotheses, and in step (2) the evidence indices from each study are fused into a single discriminatory measure representing the overall evidence against each null hypothesis. IRT has several distinct advantages. First, the IRT framework guarantees an overall FDR control as long as the individual studies control the FDR at their desired levels. This FDR control holds under arbitrary dependence between the fused evidence indices from step (2). See Section 2 for more details. Second, IRT is a powerful alternative for fusing inferences in meta-analytic settings where some studies report $p$-values while the rest reveal only the rejections under a pre-specified FDR level. Tang et al. (2014) discuss $p$-value imputation techniques in this setting assuming that all participating studies are independent. In contrast, IRT synthesizes inferences setting even when the studies are dependent. Section 3 presents this discussion. Third, IRT is extremely simple to implement and is broadly applicable without any model assumptions. This particular aspect is especially appealing because IRT synthesizes inferences from diverse studies irrespective of the underlying multiple testing algorithms employed by the studies.
The data and R-code used in this article are available at https://github.com/trambakbanerjee/IRT.
@@ 2 IRT: integrative FDR control using binary decision sequences
2.1 Notations and problem setup
We first introduce some notations and formally define the problem setup. We then introduce the three steps of the IRT framework: evidence construction, evidence aggregation, and FDR control.
In the sequel, let $\mathbb{I}(\cdot)$ denote the indicator function that returns 1 if the condition is met and 0 otherwise, denote $\|\mathbf{w}\|_{p}$ as the $\ell_{p}$-norm of vector $\mathbf{w}$, $\bm{I}_{d}$ will denote the $d\times d$ identity matrix, $\mathcal{N}_{d}(\bm{\mu},\bm{\Sigma})$ will represent the $d$-dimensional Gaussian distribution with mean vector $\bm{\mu}$ and positive definite covariance matrix $\bm{\Sigma}$, a $d$-dimensional column vector with all elements equal to a real constant $a$ will be denoted by $\bm{a}_{d}$, $d]=\{1,\ldots,d\}$ and the cardinality of a set of positive integers $\mathcal{I}$ will be denoted by $|\mathcal{I}|$.
We consider the setting of meta-analysis involving $d$ studies. For each study $j\in[d]$, a set of $m_{j}$ hypotheses, denoted by the index set $\mathcal{M}_{j}$, is tested. Let $\mathcal{M}=\bigcup_{j=1}^{d}\mathcal{M}_{j}$ denote the set of all unique hypotheses across all studies, with cardinality $|\mathcal{M}|=m$. The null hypothesis corresponding to index $i$ is denoted as $H_{0i}$. Let $\mathcal{H}_{0}=\{i\in\mathcal{M}:H_{0i}\text{ is true}\}$ be the set of true null hypotheses and $\mathcal{H}_{0j}=\{H_{0i}:i\in\mathcal{M}_{j}\cap\mathcal{H}_{0}\}$ be the set of true null hypotheses tested by study $j$. Denote $\theta_{i}=\mathbb{I}(i\notin\mathcal{H}_{0})$ the true underlying state of hypothesis $H_{0i}$. For each hypothesis $i\in\mathcal{M}_{j}$, study $j$ makes a binary decision, $\delta_{ij}\in\{0,1\}$, where $\delta_{ij}=1$ signifies a rejection of $H_{0i}$. The collection of decisions for study $j$ is represented by the vector $\bm{\delta}_{j}=(\delta_{1j},\cdots,\delta_{m_{j}j})\in\{0,1\}^{m_{j}}$. Denote $\|\bm{\delta}_{j}\|_{0}=\sum_{i\in\mathcal{M}_{j}}\delta_{ij}$ the total number of rejections made by study $j$.
A selection error, or false positive, occurs if study $j$ asserts that $H_{0i}$ is false when it is in fact true. A primary goal in multiple testing is to control the False Discovery Rate (FDR, Benjamini and Hochberg, 1995), defined as the expected proportion of false positives among all selected hypotheses. Formally, $\text{FDR}(\bm{\delta}_{j})=\mathbb{E}\left[\text{FDP}(\bm{\delta}_{j})\right]~\text{where FDP}(\bm{\delta}_{j})={\sum_{i\in\mathcal{M}_{j}}(1-\theta_{i})\delta_{ij}}/{\max\{\|\bm{\delta}_{j}\|_{0},1\}}$. The power of a testing procedure is measured by the expected proportion of true positives detected (ETP) where, $\mbox{ETP}(\bm{\delta}_{j})=\mathbb{E}[\sum_{i\in\mathcal{M}_{j}}\theta_{i}\delta_{ij}/\max\{\sum_{i\in\mathcal{M}_{j}}\theta_{i},1\}]$.
For each study $j$, we assume the triplet $\mathcal{D}_{j}=\{\bm{\delta}_{j},\alpha_{j},\mathcal{M}_{j}\}$ is available. Here, $\alpha_{j}\in(0,1)$ is the pre-specified FDR level for which the guarantee $\text{FDR}(\bm{\delta}_{j})\leq\alpha_{j}$ holds. A crucial element of our setting is that we do not always have access to the original test statistics or $p$-values that produced the decisions $\bm{\delta}_{j}$. Our objective is to synthesize the evidence from the collection of triplets $\{\mathcal{D}_{j}:j\in[d]\}$ to produce a new set of rejections for the hypotheses in $\mathcal{M}$, while controlling the overall FDR at a user-specified level $\alpha$.
2.2 The IRT framework